OpenStudy (anonymous):

What is the solution to the equation? a + (–0.4) = –3

TheSmartOne (thesmartone):

a+(-b)=-c a-b=-c a=-c+b a=b-c In this case a is a b is .4 and c is 3

TheSmartOne (thesmartone):

@Evil-cow-8

OpenStudy (anonymous):

Yes?

TheSmartOne (thesmartone):

what have you gotten as the answer...

TheSmartOne (thesmartone):

nevermind.

OpenStudy (anonymous):

3?